# Temporal Overlapping Prediction: A Self-supervised Pre-training Method for LiDAR Moving Object Segmentation

### v1

@@ 3.2 Temporal Overlapping Pre-processing
As computing temporal overlapping during pre-training can be inefficient, we pre-process the LiDAR scans $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$ to produce temporal overlapping points $\mathbf{O}$ and their occupancy states $\mathbf{S}$. We first identify coplanar beam pairs from $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$, as coplanarity is a necessary condition for beams to have overlapping observations, as illustrated in Fig. 2(a). Then, we compute the intersection of coplanar beam pairs and sample temporal overlapping points. Depending on the spatial angle between two beams, we split the computation and sampling into two scenarios, shown in Fig. 2(b) and Fig. 2(c), respectively. We take an example beam pair for a clearer explanation: the $i^{th}$ beam at time $t_{0}$ and the $j^{th}$ beam at time $t$ ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), emitted from ($\mathbf{a}^{t_{0}}$, $\mathbf{a}^{t}$) and hitting ($\mathbf{p}_{i}^{t_{0}}$, $\mathbf{p}_{j}^{t}$). When a beam pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$) generates the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$, we record the $(i,j,k)$ correspondence for subsequent usage in Sec. 3.3.
@@
$\mathbf{n}_{i}^{t}=\mathbf{d}_{i}^{t_{0}}\times\frac{\mathbf{a}^{t}}{||\mathbf{a}^{t}||}.$ (3)
The space angle between beam $\mathbf{d}_{j}^{t}$ and the plane is:
@@
Scenario-1 ($\alpha_{i,j}>\theta_{\text{dvg}}$). This scenario describes most overlapping cases where the spatial angle between two beams exceeds the beam divergence angle, as shown in Fig. 2(b). The beam intersection is small, thus being considered as a point $\mathbf{q}_{i,j}$, which is the intersection of the two beams’ centerlines. The $\mathbf{q}_{i,j}$ is solved from the parameter equations of the two beams:
@@
where $\mathbf{m}_{i,j}=\mathbf{d}_{i}^{t_{0}}\times\mathbf{d}_{j}^{t}$ is the unit common perpendicular line of the two beams. In this scenario, the intersection point $\mathbf{q}_{i,j}$ is selected as the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$.
@@
Scenario-2 ($\alpha_{i,j}\leq\theta_{\text{dvg}}$). When the spatial angle between two beams is less than the beam divergence, the intersection of the two beams can no longer be considered as a point but an area, as shown in Fig. 2(c). In this scenario, we compute the intersection segments (purple dashed lines) on the centerlines for subsequent sampling. Firstly, we obtain the endpoint $\mathbf{q}_{i,j}$ from the beam intersection similar to Scenario-1. Then, we approximate the starting point $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ as the projection from a point $\mathbf{b}$ to the two beams. Here, $\mathbf{b}$ is the crossed point of the boundaries of the two beams. Given that the spatial angle $\alpha_{i,j}$ is small, we make two approximations to compute the starting points: (1) We approximate that the two intersection segments share the same length. (2) We approximate that the distance between the two starting points $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ is equal to the sum of the beam radius at the two points. The detailed formulas for computing $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ are in the appendix.
@@
Since the overlap between two beams is an area, we sample 5 temporal overlapping points inside the intersection segment, as shown in Fig. 2(c) right: $\mathbf{o}_{k_{1}}$, the beam hit point $\mathbf{p}_{i}^{t_{0}}$; $\mathbf{o}_{k_{2}}$, the projection of $\mathbf{p}_{j}^{t}$ on $\mathbf{d}_{i}^{t_{0}}$; $\mathbf{o}_{k_{3}}$, the center of ($\mathbf{o}_{k_{1}}$, $\mathbf{o}_{k_{2}}$); $\mathbf{o}_{k_{4}}$, the center of ($\mathbf{o}_{k_{1}}$, $\mathbf{q}_{i,j}$); $\mathbf{o}_{k_{5}}$, the center of ($\mathbf{o}_{k_{2}}$, $\mathbf{q}_{i,j}$). Their occupancy states at timestamp $t$ are given by the states of their projection points on $\mathbf{d}_{j}^{t}$.
LiDAR Occupancy Measurement Model. The LiDAR occupancy measurement model [1, 43] partitions the space between the sensor origin to infinity along the beam direction into three areas with different occupancy states, as shown in Fig. 3. For the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$ generated from the overlap beams pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), the confidence level $w_{\text{conf}}$ of the measured occupancy state of $\mathbf{o}_{k}$ is 100% when $\mathbf{o}_{k}$ lies inside the physically traversed areas of the beam (blue area in Fig. 3). It then exponentially decays after the hit-point $\mathbf{p}$, as described below:
$w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t})=\begin{cases}1,\ \text{if}\ 0\leq||\mathbf{o}_{k}||\leq||\mathbf{p}_{j}^{t}||,\\ e^{{-(||\mathbf{o}_{k}||-||\mathbf{p}_{j}^{t}||)}},\ \text{if}\ ||\mathbf{o}_{k}||>||\mathbf{p}_{j}^{t}||.\\ \end{cases}$ (6)
The occupancy state of $\mathbf{o}_{k}$ is denoted as $\mathbf{s}_{k}$, which is described as:
$\mathbf{s}_{k}=\begin{cases}[1,0,0],\ \text{if}\ 0\leq||\mathbf{o}_{k}||<||\mathbf{p}_{j}^{t}||,\\ [0,1,0],\ \text{if}\ ||\mathbf{o}_{k}||=||\mathbf{p}_{j}^{t}||\ \text{or}\ \lambda_{\text{occ}}\leq w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t}),\\ [0,0,1],\ \text{if}\ w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t})<\lambda_{\text{occ}},\\ \end{cases}$ (7)
@@
where $\mathbf{s}_{k}$ is a one-hot vector indicating “free”, “occupied” and “unknown” occupancy states, and $\lambda_{\text{occ}}$ is the confidence level threshold of the occupied state.
@@ 3.3 Training Objectives
For self-supervised pre-training, the primary objective is temporal overlapping prediction, which involves predicting $\mathbf{s}_{k}$ from the $i^{th}$ current point’s feature $\mathbf{f}_{i}$ and the positional information of $\mathbf{o}_{k}$ using a shallow MLP decoder $h(\cdot)$:
@@
where $\mathbf{u}_{k}$ is the predicted probabilities of three occupancy states, and $f(\mathbf{o}_{k})\in\mathbb{R}^{d}$ denotes the sine-cosine 4D positional encoding vector as described in [49]. Temporal overlapping prediction is described as a Cross-Entropy loss:
@@
where $\mathbf{w}_{s}$ is a constant weight vector for each occupancy state, and the $(i,j,k)$ correspondence is introduced in Sec. 3.2.
Additionally, we reconstruct the current occupancy using the input scans, which serves as a regularization term in the total loss. We randomly select $M^{\prime}$ occupied and free points on each current beam. For the $k^{th}$ point $\mathbf{c}_{k}$ sampled on the $i^{th}$ current beam, it includes the spatial position and timestamp $t_{0}$. The ground truth occupancy label is $\mathbf{s}_{k}$, and the predicted probability is $\mathbf{v}_{k}$. The reconstruction objective is also specified as a Cross-Entropy loss:
@@
$\mathcal{L}_{\text{recon}}$ $=-\frac{1}{NM^{\prime}}\sum\limits_{k=1}^{NM^{\prime}}\mathbf{w}_{s}\cdot\mathbf{s}_{k}\cdot\log{(\mathbf{v}_{k})}.$ (11)
@@ 4 Evaluation Metrics
In this section, we propose a mean object-level metric denoted as $\text{mIoU}_{\text{obj}}$ to tackle the biased weighting issue in the conventional IoU metric. Previous methods [9, 33, 34, 50] have relied solely on IoU as the evaluation metric. However, IoU implicitly gives larger weight to objects with more scanned points due to its point-wise nature. Objects of varying geometric sizes and distances to the LiDAR may result in different numbers of scanned points. Nevertheless, these objects are equally important in practical MOS applications and should not be weighted based on the number of scanned points. To further illustrate the impact of this bias on MOS performance evaluation, we conduct a statistical analysis: we sort all moving objects in the nuScenes train-val split based on the number of scanned points, and then calculate the cumulative distribution of moving objects and scanned points. The results indicate that the top 75.22% of objects with the fewest scanned points only contribute to 15.49% of the total points, while 90.06% of objects contribute to 37.02% of the points, revealing that a few objects with more points heavily influence the conventional IoU metric. The cumulative distribution is provided in the appendix.
The proposed metric $\text{mIoU}_{\text{obj}}$ calculates the IoU for each moving object and computes the average, ensuring equal weighting for each moving object:
$\text{mIoU}_{\text{obj}}=\frac{1}{m}\sum\limits_{i=1}^{m}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}},$ (12)
where $m$ denotes the number of moving objects, $\text{TP}_{i}$ and $\text{FP}_{i}$ indicate the true positive and false positive classifications of the points belonging to $i^{th}$ moving object. False negatives are not involved in the $\text{mIoU}_{\text{obj}}$ as all points are classified as the moving (positive) class.
Additionally, in the nuScenes train-val split, ego vehicles’ moving points account for 76.72% of all moving points, which can heavily influence the IoU. However, in autonomous driving scenarios, we already know the movement of the ego vehicle and do not need to predict it. Therefore, we utilize a refined IoU metric called $\text{IoU}_{\text{w/o}}$, calculated as described in [9, 33, 34] but excluding the scanned ego vehicle points.
For a more precise and robust evaluation of the MOS performance, we use both metrics: $\text{mIoU}_{\text{obj}}$ assesses performance on all moving objects, while $\text{IoU}_{\text{w/o}}$ reflects overall performance, incorporating false negatives.
@@ 5.1 Experiments Setup
Datasets. We pretrain the encoder on the nuScenes dataset [6], which utilizes a 32-beam Velodyne LiDAR for data collection. The beam divergence angle $\theta_{\text{dvg}}$ specified by the manufacturer is 0.003 rad. We perform few-shot experiments on the nuScenes dataset. Previous methods [34, 46] convert nuScenes object attributes into MOS labels. However, these attributes do not precisely describe the object’s motion state. Moreover, some objects lack attributes and will be overlooked during training or evaluation. To overcome this limitation, we refine MOS labels by incorporating object attributes and speed. Speed is computed from object bounding boxes and is used to classify motion states by static and moving speed thresholds. We introduce a speed range between these thresholds to categorize objects with unclear motion states into an unknown class, such as cars that are just starting or hesitant pedestrians. More details are available in the appendix. We also conduct few-shot cross-dataset transfer experiments on the SemanticKITTI [3] MOS benchmark [9], which employs a 64-beam Velodyne LiDAR for data collection.
Implementation Details. In our setup, we have $n=6$, where $\mathcal{P}_{\text{adj}}$ includes 6 past scans and 6 future scans, and the input sequence $\mathcal{P}_{\text{in}}$ comprises 6 scans. All scans are sampled at 2Hz. The confidence cutoff threshold for the occupied state is $\lambda_{\text{occ}}=0.9$. We apply data augmentation techniques from [37, 33] and discard points outside a bounded area based on [23, 1]. Our encoder is the sparse 4D UNet backbone, commonly used in MOS methods [33, 34]. The quantization size for the sparse backbone is 0.1m, and the feature dimension $d=128$. For the temporal overlapping prediction objective, all occupied class samples are included due to their scarcity, while 5 times the number of occupied samples are randomly selected for the free class and the same amount for the unknown class in each training iteration. As for the current occupancy reconstruction objective, we select 5 occupied points and 25 free points on each current beam ($M^{\prime}=30$). The class weight is $\mathbf{w}_{s}=[1,5,1]$ for free, occupied, and unknown class. We use a fixed random seed for all experiments to ensure reproducibility. Detailed optimizer settings are in the appendix.
Baselines. We use the SOTA supervised MOS method 4DMOS [33] as the training-from-scratch baseline. Additionally, we implement the SOTA LiDAR observation forecasting methods 4Docc [23] and UnO [1], along with the occupancy-based 3D method ALSO [4], as the self-supervised pre-training baselines. All baselines utilize the same sparse 4D UNet [12, 33] encoder with their respective original decoders.

### v2

@@ 3.2 Temporal Overlapping Pre-processing
As computing temporal overlapping points during pre-training can be inefficient, we pre-process the LiDAR scans $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$ to produce temporal overlapping points $\mathbf{O}$ and their occupancy states $\mathbf{S}$. We first identify coplanar beam pairs from $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$, as coplanarity is a necessary condition for beams to overlap, as illustrated in Fig. 3(a). Then, we compute the intersection of coplanar beam pairs and sample temporal overlapping points. We handle the computation and sampling in two distinct scenarios based on the spatial angle between two beams, as shown in Fig. 3(b) and Fig. 3(c), respectively. For clarity, we consider a beam pair: the $i^{th}$ beam at time $t_{0}$ and the $j^{th}$ beam at time $t$ ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), emitted from ($\mathbf{a}^{t_{0}}$, $\mathbf{a}^{t}$) and hitting ($\mathbf{p}_{i}^{t_{0}}$, $\mathbf{p}_{j}^{t}$). When a beam pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$) generates the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$, we record the $(i,j,k)$ correspondence for subsequent usage in Sec. 3.3.
@@
[…]
The spatial angle between beam $\mathbf{d}_{j}^{t}$ and the plane is:
@@
Scenario-1 ($\alpha_{i,j}>\theta_{\text{dvg}}$). This scenario describes most overlapping cases where the spatial angle between two beams exceeds the beam divergence angle, as shown in Fig. 3(b). The beam intersection is small and is thus considered a point $\mathbf{q}_{i,j}$, which is the intersection of the two beams’ centerlines. $\mathbf{q}_{i,j}$ is calculated from the parameter equations of the two beams:
@@
where $\mathbf{m}_{i,j}=\mathbf{d}_{i}^{t_{0}}\times\mathbf{d}_{j}^{t}$ is a vector along the common perpendicular to the two beams. In this scenario, the intersection point $\mathbf{q}_{i,j}$ is selected as the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$.
@@
Scenario-2 ($\alpha_{i,j}\leq\theta_{\text{dvg}}$). When the spatial angle between two beams is less than or equal to the beam divergence, the intersection of the two beams can no longer be considered a point but rather a region, as shown in Fig. 3(c). In this scenario, we compute the intersection segments (purple dashed lines) on the centerlines for subsequent sampling. Firstly, we obtain the endpoint $\mathbf{q}_{i,j}$ from the beam intersection similar to Scenario-1. Then, we approximate the starting point $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ as the projection from a point $\mathbf{b}$ onto the two beams. Here, $\mathbf{b}$ is the crossing point of beam boundaries. Given that the spatial angle $\alpha_{i,j}$ is small, we make two approximations to compute the starting points: (1) We approximate that the two intersection segments share the same length. (2) We approximate that the distance between the two starting points $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ is equal to the sum of the beam radii at the two points. The detailed formulas for computing $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ are provided in the appendix.
@@
Since the overlap between two beams is a region, we sample 5 temporal overlapping points within the intersection segment, as shown in Fig. 3(c)-right: $\mathbf{o}_{k_{1}}$, the beam hit point $\mathbf{p}_{i}^{t_{0}}$; $\mathbf{o}_{k_{2}}$, the projection of $\mathbf{p}_{j}^{t}$ on $\mathbf{d}_{i}^{t_{0}}$; $\mathbf{o}_{k_{3}}$, the midpoint of ($\mathbf{o}_{k_{1}}$, $\mathbf{o}_{k_{2}}$); $\mathbf{o}_{k_{4}}$, the midpoint of ($\mathbf{o}_{k_{1}}$, $\mathbf{q}_{i,j}$); $\mathbf{o}_{k_{5}}$, the midpoint of ($\mathbf{o}_{k_{2}}$, $\mathbf{q}_{i,j}$). The occupancy state for each of these sampled points at timestamp $t$ is determined by projecting it onto the adjacent beam $\mathbf{d}_{j}^{t}$ and evaluating the corresponding occupancy.
LiDAR Occupancy Measurement. As shown in Fig. 2, the LiDAR occupancy measurement [1, 43] partitions the space along each beam into three states. First, the space from the sensor origin to the reported point is considered free with maximum confidence. Beyond the reported point, an exponential confidence decay defines two subsequent states: the space is occupied while the confidence exceeds a predefined threshold, and is considered unknown where the confidence drops below the threshold.
Given the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$ generated from the overlapping beam pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), the confidence level $w_{\text{conf}}$ of $\mathbf{o}_{k}$ is formulated as:
[…]
The occupancy state of $\mathbf{o}_{k}$ at time $t$ is denoted as $\mathbf{s}_{k}$:
[…]
@@
where $\mathbf{s}_{k}$ is a one-hot vector indicating one of three occupancy states: free, occupied, or unknown. $\lambda_{\text{occ}}$ is the predefined confidence threshold for the occupied state.
@@ 3.3 Training Objectives
Our primary pre-training objective, temporal overlapping prediction, is to predict the occupancy state $\mathbf{s}_{k}$ of $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$. We predict the occupancy based on the positional encoding of $\mathbf{o}_{k}$ and the $i^{th}$ current LiDAR point’s feature $\mathbf{f}_{i}$ by a shallow MLP decoder $h(\cdot)$:
@@
where $\mathbf{u}_{k}$ is the predicted probability over the three occupancy states, and $f(\mathbf{o}_{k})\in\mathbb{R}^{d}$ denotes the sine-cosine 4D positional encoding vector as described in [49]. Temporal overlapping prediction is formulated as a Cross-Entropy loss:
@@
where $\mathbf{w}_{s}$ is a constant weight vector for each occupancy state, and the $(i,j,k)$ correspondence was introduced in Sec. 3.2.
Additionally, we reconstruct the current occupancy using the input scans, which serves as a regularization term in the overall loss. We randomly sample $M^{\prime}$ occupied and free points along each current beam. For the $k^{th}$ point $\mathbf{c}_{k}$ sampled on the $i^{th}$ current beam, it consists of a spatial position and timestamp $t_{0}$. The ground truth occupancy label is $\mathbf{s}_{k}$, and the predicted probability is $\mathbf{v}_{k}$. The reconstruction objective is also formulated as a Cross-Entropy loss:
@@
$\mathcal{L}_{\text{recon}}$ $=-\frac{1}{N_{t_{0}}M^{\prime}}\sum\limits_{i=1}^{N_{t_{0}}}\sum\limits_{k=1}^{M^{\prime}}\mathbf{w}_{s}\cdot\mathbf{s}_{k}\cdot\log{(\mathbf{v}_{k})}.$ (11)
@@ 4 Evaluation Metrics
We introduce a robust evaluation for MOS, designed to overcome two critical biases in the conventional IoU metric [9, 33, 34, 50]: the over-weighting of objects with more points, and the inflation caused by the ego vehicle’s points.
A primary issue with the conventional IoU is its inherent bias towards objects with more points, leading to the neglect of smaller or distant objects. This poses significant safety risks, as all dynamic objects are equally critical for safe operation, regardless of their point count. For instance, a high-speed vehicle might fail to detect a distant car, or a UAV could miss a fast-approaching object. To quantify this imbalance, our statistical analysis of the nuScenes dataset reveals that the smallest 75.22% of objects (by point count) account for only 15.49% of the total points (see appendix). This confirms that a few large objects heavily dominate the final IoU score. The second bias is the inflation of IoU. In practice, the ego vehicle’s state is known via onboard sensors, and is not part of the perception challenge. Including these easy-to-predict points (which account for 76.72% of all moving points in the nuScenes dataset) severely inflates the metric, masking the model’s actual performance on external dynamic objects.
To address these limitations, we use two complementary metrics. First, to counteract the object-size bias, we use an object-level metric $\text{Recall}_{\text{obj}}$, which ensures an equal weight of each moving object by averaging their individual recalls. It is computed as:
$\text{Recall}_{\text{obj}}=\frac{1}{m}\sum\limits_{i=1}^{m}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}},$ (12)
where $m$ is the number of moving objects, while $\text{TP}_{i}$ and $\text{FN}_{i}$ are the number of true positive and false negative points, respectively, within the ground truth bounding box of the $i^{th}$ object. Second, to resolve the ego-vehicle bias, we simply exclude all ego points from the IoU calculation and rename the metric $\text{IoU}_{\text{w/o}}$. This provides a more focused assessment of the model’s performance on external dynamic objects. Notably, for the SemanticKITTI benchmark, $\text{IoU}_{\text{w/o}}$ is equivalent to the conventional IoU, as the benchmark already omits ego points by classifying them as unlabeled. To demonstrate that the conventional IoU metric fails to effectively evaluate perception of external dynamic objects when numerous ego-vehicle points are included, we provide supplementary results on nuScenes using conventional IoU in the appendix.
In summary, our dual-metric evaluation provides a holistic and unbiased assessment. The two metrics are complementary: $\text{Recall}_{\text{obj}}$ ensures that the detection completeness of each object is fairly evaluated, preventing smaller ones from being overlooked, while $\text{IoU}_{\text{w/o}}$ penalizes false positive predictions in the surrounding scene.
@@ 5.1 Experiments Setup
Datasets. We pretrain the encoder on the nuScenes dataset [6], which utilizes a 32-beam Velodyne LiDAR for data collection. The beam divergence angle $\theta_{\text{dvg}}$ specified by the manufacturer is 0.003 rad. We perform few-shot experiments on the nuScenes dataset. Previous methods [34, 46] directly convert nuScenes object attributes into MOS labels. However, these attributes do not precisely describe the object’s motion state. Moreover, some objects lack attributes and are thus overlooked during training or evaluation. To overcome this limitation, we refine MOS labels by incorporating object attributes and speed. We compute object speed from bounding box annotations and use it to classify motion states based on predefined static and moving thresholds. We introduce a speed range between these thresholds to categorize objects with ambiguous motion states into an unknown class, such as cars that are just starting or hesitant pedestrians. More details are available in the appendix. We also conduct few-shot cross-dataset transfer experiments on the SemanticKITTI [3] MOS benchmark [9], which employs a 64-beam Velodyne LiDAR for data collection.
Implementation Details. In our setup, we set $n=6$, where $\mathcal{P}_{\text{adj}}$ includes 6 past scans and 6 future scans, and the input sequence $\mathcal{P}_{\text{in}}$ comprises 6 scans. All scans are sampled at 2Hz. The confidence cutoff threshold for the occupied state is $\lambda_{\text{occ}}=0.9$. We apply data augmentation techniques from [37, 33] and discard points outside a bounded area based on [23, 1]. Our encoder is a sparse 4D UNet. Following previous works [33, 34], we initialize all point features to a constant. The quantization size for the sparse backbone is 0.1 m, and the feature dimension $d=128$. For our temporal overlapping prediction objective, we address class imbalance due to the scarcity of occupied samples. Specifically, in each training iteration, we use all available occupied samples, while randomly sampling five times as many points for the free class and the same number of points for the unknown class. As for the current occupancy reconstruction objective, we sample 5 occupied points and 25 free points on each current beam ($M^{\prime}=30$). The class weight is $\mathbf{w}_{s}=[1,5,1]$ for free, occupied, and unknown classes. We use a fixed random seed for all experiments to ensure reproducibility. Detailed optimizer settings are in the appendix.
Baselines. We benchmark our method against two classes of baselines: (1) a training-from-scratch baseline using the sparse 4D UNet from [33, 34, 50], and (2) self-supervised pre-training baselines, including the LiDAR forecasting methods 4DOcc [23] and UnO [1], and the occupancy-based 3D pre-training method ALSO [4]. To ensure a fair comparison, we replace their native encoders with the sparse 4D UNet while retaining their original decoders.